The market for Brussels Sprouts in California is represented by the equations below. Q is measured in thousands of bushels.

P = 9 – Qd        P = 2Qs

PLEASE SHOW STEPS :)

  1. What is the equilibrium price and quantity of brussels sprouts (in thousands of bushels)?
  2. Graph the two equations, labeling the axes (with units) and equilibrium price and quantity.

Imagine the government of California is considering placing a price floor of $8 per bushel on brussels sprouts.

  1. Is this price floor binding? How do you know?
  2. What are the quantities supplied and demanded with the price floor? [Hint: you can find Qs and Qd by plugging the new price into the supply and demand equations]
  3. What type of disequilibrium situation is created, and what is its magnitude? How many thousand bushels of brussels sprouts will be sold under this policy?

Now let’s look at how Consumer, Producer, and Total Surplus are affected by this policy.

  1. Without calculating the values, state whether each of CS, PS, and TS will increase, decrease, or stay the same. [Hint: Looking at the areas on the graph can help to determine the changes.]
  2. Which parties are made better off by the price floor? Who is worse off?
  3. What is the dollar value of the deadweight loss caused by this policy? [Hint: You can either use the graph to find the area of the triangle, or calculate the change in total surplus.]
Answer #1

a) Set QD=QS

9-Q=2Q

9 = 3Q

Q = 9/3 =3

P = 9-3 =6

b) When QD=0, P=9 and When QS=0, P =0

c) Yes, this is a binding price floor as a binding price floor is imposed above the equilibrium price

d) When P=8,

QD: 1

QS: 4

e) As quantity demanded is lower than the quantity supplied so, there exists a surplus in the market

Surplus = QS-QD = 4-1 =3 units

The market will have a deadweight loss as equilibrium quantity will fall equal to the quantity demanded and quantity sold = quantity demanded = 1

As per Chegg guidelines, more than four parts are answered.